OpenStudy (anonymous):

this from a calc text: integral of sqrt(x-1)/x wolfram spits it out but gives no indication as to what substitution was made

OpenStudy (anonymous):

\[\int (x-1)^{0.5}x^{-1} dx\] let u=sqrt(x-1)

OpenStudy (anonymous):

\[u=\sqrt(x-1), x=u^2+1, dx=2\sqrt{x-1}du, dx=2u\] \[\int (x−1)^{0.5}x^{−1}dx= \int\frac{u}{u^2+1} 2udu\]\[= 2\int \frac{u^2}{u^2+1} du\]\[=2\int1-\frac{1}{u^2+1}du\]\[=2u-2tan^{-1}{u}+c\]\[=2\sqrt{x-1}-2tan^{-1}{\sqrt{x-1}}+c\]

OpenStudy (anonymous):

thank you very much Zed. i was going along that very same path to solution-i failed to make that algebraic operation where you broke the integrand into 2 parts

myininaya (myininaya):

u^2=u^2+1-1 :)

OpenStudy (anonymous):

yeah long divisions a pain :) here's a website that has the inverse trig integrals http://www.math.com/tables/integrals/tableof.htm

OpenStudy (anonymous):

ok

myininaya (myininaya):

sometimes doing the long division is easier though

myininaya (myininaya):

but definitely not in this case gj by the way zed

OpenStudy (anonymous):

ok, so you divided u^2 by (u^2+1) to change the expression to 2 parts. now I see thanx again